A charge -Q is between a charge Q and a charge 2Q separated by a distance d. How far from the charge Q is the negative charge if the net force on the negative charge is 0?
Answer #1

As given in question, charge "-Q" is in between charge "+Q" and charge "+2Q".

We assume that charge "-Q" is "x" distance apart from charge "+Q". so, "-Q" will be (d-x) distant from charge "+2Q".

Force on "-Q" by charge "Q" will be ; F1 = 1/(4π *ε) * (Q*Q)/x2

Force on "-Q" by charge "2Q" will be ; F2 = 1/(4π *ε) * (2Q*Q)/(d-x)2

As net force is zero, F1 and F2 must be equal (We already know that it is in opposite direction)

1/(4π *ε) * (Q*Q)/x2  = 1/(4π *ε) * (2Q*Q)/(d-x)2

1/x2 = 2/ (d-x)2

[(d-x)/x]2 = 2

(d-x)/x = √2

(d/x) -1 = √2

d/x = √2 + 1

x = d / (√2 +1)